Circular No. 6055

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS R. M. Hjellming, National Radio Astronomy Observatory, reports:

“Very Large Array radio observations of the field of GRO J1655–40 on Aug. 12.18 UT at frequencies of 1.4, 4.9, and 14.9 GHz show the presence of a very strong radio source located at α = 16^{h}54^{m}00^{s}.20, δ = −39^{o}50′43′′.6 (equinox 2000.0; ± 0′′.5), with flux densities of 1.04, 1.13, and 0.76 Jy, respectively (± 5 percent). This radio source was unresolved (≤ 0′′.5), and — with a location only 3′′ from the position of the optical counterpart reported by Bailyn et al. (IAUC 6050, with a position uncertainty of 2′′) — is coincident with both that optical counterpart and the variable radio source observed by Campbell-Wilson and Hunstead (IAUC 6052). The Molonglo Synthesis Telescope observations at 843 MHz, showed flux densities of 0.37, 0.77, and 0.9 Jy on Aug. 6.60, 11.59, and 12.33, respectively — so the rising radio counterpart of the x-ray nova had a non-thermal synchrotron radio spectrum on Aug. 12 with a peak near 5 GHz. This is the strongest optically-thick radio emission ever seen for a soft x-ray transient, and the strongest seen in any state two weeks after outburst. It may be similar to the long-lived, optically-thick radio source associated with the transient GS 2023+338 (V404 Cyg; Han and Hjellming 1992, Ap.J. 400, 304).”

D. Campbell-Wilson and R. Hunstead, University of Sydney, report the results of further radio observations of the field of GRO J1655–40, using the Molonglo Observatory Synthesis Telescope at 843 MHz: “Following our previous report (IAUC 6052), the radio emission has increased spectacularly, reaching 4.2 Jy (± 5 percent) at Aug. 14.58 UT and 5.5 Jy (± 5 percent) at Aug. 15.19. If we assume an exponential increase from Aug. 11 to 15, the implied time constant is 1.8 days. However, the overall rate of increase from Aug. 6 to 15 appears to be faster than exponential, even allowing for a possible quiescent level. Observations at other frequencies are essential for monitoring the spectrum of the emission and its evolution with time.”

COMET MACHHOLZ (1994o) H. Mikuž, Ljubljana, Slovenia, reports that a CCD image with a 0.20-

m f/2 Baker-Schmidt camera yields V = 11.0 on Aug. 15.97 UT. A 1min exposure on Aug. 16.009 shows a prominent, slightly curved jet of gas extending ∼ 3′ out of the comet nucleus in p.a. 350^{o}.

1994 August 16 Daniel W. E. Green